# Very Basic Set theory

### v5

@@
Very Basic Set theory
Doeko H. Homan
(Date: February 5, 2025)
George Boolos describes the naive conception of set ”Any predicate has an extension” (Boolos [1971] pp.215-231). Then Russell’s paradox is used to prove the inconsistency of naive set theory. Russell’s paradox reads
We call a set which contains itself an ‘individual’.
In this article we show Russell’s paradox is logically equivalent to:
For every set $s$, an individual is contained in $s$ or a set that is not an individual
is not contained in $s$.
Very basic set theory is a set theory with individuals that has its philosophical foundation in Ludwig Wittgenstein’s Tractatus Logico-Philosophicus.
@@ Introduction
In (Boolos [1971] pp.216-217) ”individuals (= non-sets)”.
@@
Thus ‘$x$ $\varepsilon$ $x$ or not’ is definite; the possibility ‘$x$ $\varepsilon$ $x$’ is not excluded by our axioms.
@@
Then for every $s$ it is the case $s\in s$ or $\notin s$. We prove Russell’s paradox
Assume exists $s$ such that for every $u$ ($u\in s$ if and only if $u\notin u$).
The formula applies to every $u$ thus we can choose $s$ therefore
exists $s$ such that ($s\in s$ if and only if $s\notin s$).
The question ‘is $s$ contained in $s$’ or ‘is $s$ not contained in $s$’ is definite but there is
no answer: If $s\in s$ then $s\notin s$ and vice versa. Therefore $s$ is nonexisting. $\Box$
@@ 2. What is a set?
At an early age you develop the mental picture of ‘what is a set’. You belong to a family or a clan, to the inhabitants of a village or a particular region. Then experience shows there are things that constitute a set. The members of a set are the constituents of that set. Therefore set theory describes
@@ 3. Logic and definiteness
If it is not the case that both $\Phi$ and $\neg\Phi$, that is formula $\neg(\Phi\wedge\neg\Phi)$, then ‘$\Phi$ is definite’. For every $s$ and for every $t$ it is not the case that both $s\in t$ and $\neg(s\in t)$ thus formula $s\in t$ is definite.
@@
If $\Phi$ is definite then $\forall s(\Phi)$ and $\exists s(\Phi)$ are definite otherwise $(\forall s(\Phi)\wedge\neg\forall s(\Phi))$ or $(\exists s(\Phi)\wedge\neg\exists s(\Phi))$ thus $\exists s(\Phi\wedge\neg\Phi)$.
@@
If $\Phi$ is definite then it is not the case that both $\Phi$ and $\neg\Phi$. Thus it is not the case that both ¬($\neg(\neg\Phi)$ and $\neg\Phi$ therefore $\neg\Phi$ is definite.
And $(\Phi\wedge\Psi)\wedge\neg(\Phi\wedge\Psi)$ is logically equivalent to $((\Phi\wedge\neg\Phi)\wedge\Psi)\vee(\Phi\wedge(\Psi\wedge\neg\Psi))$ therefore if $\Phi$ and $\Psi$ are definite then $(\Phi\wedge\Psi)$ is definite.
The logical ‘or’, ‘implies’ and ‘if and only if’ are expressed in terms of logical ‘not’ and logical ‘and’. Thus if $\Phi$ and $\Psi$ are definite then
$\neg\Phi$, $\Phi\wedge\Psi$, $\Phi\vee\Psi$, $\Phi\rightarrow\Psi$, $\Phi\leftrightarrow\Psi$, $(\Phi)$, $\forall s(\Phi)$ and $\exists s(\Phi)$ are definite.
Therefore the building up of formulas preserves definiteness.
@@ 4. Russell’s paradox
We prove formula $\neg\exists s\forall u(u\in s\leftrightarrow u\notin u)$, that is Russell’s paradox
@@
$\forall s\exists u(u\in s\leftrightarrow u\in u)$ thus $\forall s\exists u((u\in s\wedge u\in u)\vee(u\notin u\wedge u\notin u))$.
@@ 7. Axiom ‘pairs’
However, if $s\neq t$ and $(s\in t\wedge t\in s)$ then $s\notin s$ and $t\notin t$. Then exists set $\{s,t\}$ thus there is the ‘vicious membership circle’ $(s\in t\wedge t\in s)$.
Therefore we postulate ‘regularity’.
@@ 8. Axiom ‘regularity’
And it is not the case that exists $s$ such that
@@
$t\in s$, $u\in t$, $v\in u$ is nonexisting.
Thus $\forall s\forall t(s\neq t\rightarrow(s\notin t\vee t\notin s))$.
@@ 10. Natural numbers
It is straightforward to define a number sequence satisfying the Peano Axioms (PA) for natural numbers. The first axiom of PA reads: There is a first natural number. Natural numbers are also known as ‘finite ordinals’.
@@
Then $\alpha$ is a natural number. In physical reality a first number is called ‘the unit of measurement’. The classical Greek concept of number (arithmos) is ”a finite plurality composed of units, where a unit is whatever counts as one thing in the number under consideration.” (Mayberry [2000] p.21).
@@
Any pair of individuals is a first number of a natural number sequence. Therefore ‘the set of all natural numbers’ is nonexisting. However, to postulate ‘set $\omega$ of all natural numbers with first number $\alpha$’ is not excluded. Then $\bigcup\omega=\omega$. And if $\bigcup\beta=\beta$ we can define ‘transfinite ordinal numbers with first number $\beta$’ analogous to natural numbers (finite ordinal numbers) with first number $\alpha$.
@@
Thus the logic in Very basic set theory is adequate.
@@ Conclusion
Very basic set theory is a set theory that has its philosophical foundation in Ludwig Wittgenstein’s Tractatus Logico-Philosophicus. It is a set theory with individuals, the logic of definiteness is adequate and the naming of individuals is logically first. But individuals are not needed for mathematical purposes therefore individuals are banished from ‘mainstream set theory’.
@@
Set $v$ is denoted by ${\mathcal{P}}s$. Then ${\mathcal{P}}\{s\}=\{\{s\}\}$ thus the ‘number of members’ of $\{s\}$ is equal to the ‘number of members’ of ${\mathcal{P}}\{s\}$.
Cantor’s Theorem (Zermelo [1908] p.276) reads: For every set $s$,
the ‘number of members’ of $s$ is less than the ‘number of members’ of ${\mathcal{P}}s$.
Thus Cantor’s Theorem does not hold.

### v6

@@
Abstract.
The naive conception of set is described as ”Any predicate has an extension”. Then Russell’s paradox is used to prove the inconsistency of naive set theory. Russell’s paradox reads ”No set can contain all and only those sets which do not contain themselves”. We call a set which contains itself an ‘individual’. In this article we show Russell’s paradox is logically equivalent to ‘for every set $s$, an individual is contained in $s$ or a set that is not an individual is not contained in $s$’. Very basic set theory is a set theory with individuals that has its philosophical foundation in Ludwig Wittgenstein’s Tractatus Logico-Philosophicus.
@@ Introduction
In (Oliver and Smiley [2006] pp.145-146) Alex Oliver and Timothy Smiley explain
individuals are not needed therefore are prohibited in set theory.
@@
In (Boolos [1971] pp.216-217) George Boolos notes ”individuals (= non-sets)”.
@@
Thus ‘$x$ $\varepsilon$ $x$ or not’ is definite; the possibility ”$x$ $\varepsilon$ $x$” is not excluded by our axioms.
@@
Then for every set $s$ it is the case $s\in s$ or $s\notin s$, but not both. We define
  • a set is an ‘individual’ if the only constituent of that set is that set itself.
In Section 1 and Section 2 we describe individuals and sets, in Section 3 we prove logical formulas preserve definiteness. Then Russell’s paradox is proved and axioms specification, equality, pairs, regularity and union are formulated. Some important theorems are proved, and in Section 10 natural numbers are defined.
@@ 2. What is a set?
At an early age you develop the mental picture of ‘what is a set’. You belong to a family or a clan, to the inhabitants of a village or a particular region. Then experience shows there are things that constitute a set. The members of a set are the constituents of that set. Therefore Very basic set theory describes
@@ 3. Logic and definiteness
We define
  • $\Phi$ is definite’ if $\neg(\Phi\wedge\neg\Phi)$ thus it is not the case that both $\Phi$ and $\neg\Phi$.
For every $s$ and for every $t$ it is not the case that both $s\in t$ and $\neg(s\in t)$ thus $s\in t$ is definite We prove: The building up of logical formulas preserves definiteness.
If $\Phi$ is definite then $\forall s(\Phi)$ and $\exists s(\Phi)$ are definite otherwise
$(\forall s(\Phi)\wedge\neg\forall s(\Phi))$ or $(\exists s(\Phi)\wedge\neg\exists s(\Phi))$ thus $\exists s(\Phi\wedge\neg\Phi)$,
If $\Phi$ is definite then it is not the case that both $\Phi$ and $\neg\Phi$ thus
it is not the case that both ¬($\neg(\neg\Phi)$ and $\neg\Phi$ therefore $\neg\Phi$ is definite,
$(\Phi\wedge\Psi)\wedge\neg(\Phi\wedge\Psi)$ is logically equivalent to $((\Phi\wedge\neg\Phi)\wedge\Psi)\vee(\Phi\wedge(\Psi\wedge\neg\Psi))$ thus if $\Phi$ and $\Psi$ are definite then $(\Phi\wedge\Psi)$ is definite,
The connectives $\vee$, $\rightarrow$ and $\leftrightarrow$ are expressed in terms of $\neg$ and $\wedge$ thus if $\Phi$ and $\Psi$ are definite then $\Phi\vee\Psi$, $\Phi\rightarrow\Psi$ and $\Phi\leftrightarrow\Psi$ are definite.
Therefore the building up of logical formulas preserves definiteness. $\Box$
@@ 4. Russell’s paradox
We prove Russell’s paradox, that is formula $\neg\exists s\forall u(u\in s\leftrightarrow u\notin u)$
@@
$\forall s\exists u(u\in s\leftrightarrow u\in u)$ thus $\forall s\exists u((u\in s\wedge u\in u)\vee(u\notin s\wedge u\notin u))$.
@@ 7. Axiom ‘pairs’
However, if $s\neq t$ and $(s\in t\wedge t\in s)$ then $s\notin s$ and $t\notin t$. Then exists set $\{s,t\}$ thus there is the ‘vicious membership circle’ $(s\in t\wedge t\in s)$. That is contradictory to ‘the naming of individuals is logically first’.
@@ 8. Axiom ‘regularity’
Then it is not the case that exists $s$ such that
@@
$t\in s$, $u\in t$, $v\in u$ is contradictory to ‘the naming of individuals is logically first’. Then $\forall s\forall t(s\neq t\rightarrow(s\notin t\vee t\notin s))$.
@@ 10. Natural numbers
It is straightforward to define a number sequence satisfying the Peano Axioms for natural numbers. The first axiom of the Peano Axioms reads ‘there is a first natural number’ (denoted in (Mendelson [2015] p.153) by ‘00’).
@@
Then $\alpha$ is a first natural number. In physical reality a first number is called ‘the unit of measurement’. The classical Greek concept of number (arithmos) is ”a finite plurality composed of units, where a unit is whatever counts as one thing in the number under consideration.” (Mayberry [2000] p.21).
@@
Any pair of individuals is a first number of a natural number sequence. Therefore ‘the set of all natural numbers’ is nonexisting. However, to postulate ‘set $\omega$ of all natural numbers with first number $\alpha$’ is not excluded. Then $\bigcup\omega=\omega$. Therefore we can define ‘transfinite ordinal numbers with first number $\omega$’ analogous to natural numbers (the finite ordinal numbers) with first number $\alpha$.
@@
The logic in Very basic set theory is capable to deal with transfinite numbers.
@@ Conclusion
Very basic set theory is a set theory that has its philosophical foundation in Ludwig Wittgenstein’s Tractatus Logico-Philosophicus. It is a set theory with individuals, the logic is capable to deal with transfinite numbers and the naming of individuals is logically first. But individuals are not needed for mathematical purposes therefore individuals are banished from ‘mainstream set theory’.
@@
Set $v$ is denoted by ${\mathcal{P}}s$. Then ${\mathcal{P}}\{s\}=\{\{s\}\}$ thus the ‘number of members’ of $\{s\}$
is equal to the ‘number of members’ of ${\mathcal{P}}\{s\}$.
Cantor’s Theorem (Zermelo [1908] p.276) reads
for every set $s$, the ‘number of members’ of $s$ is less than the ‘number of
members’ of ${\mathcal{P}}s$.
Thus Cantor’s Theorem does not hold. $\Box$